OpenStudy (anonymous):

At a restaurant, Trevor decides to tip the server 16% of the bill. If he leaves a $4.00 tip, which proportion could be used to find how much the bill was for?

OpenStudy (anonymous):

if 16%=4$ then 1$=4% then 100%=25$

OpenStudy (anonymous):

i need a proportion

OpenStudy (anonymous):

x/100=4/16

OpenStudy (anonymous):

did you get it

OpenStudy (anonymous):

it has to be 16/100=x/4 or 16/100=4/x